Ewa needs a medal for keeping six guys relatively civilised for the last weeks, but the prospect of Jenine, Liane and Steffi coming in today has led to the mess tent being re-pitched and a general tidying up around the camp. In fairness to the guys, this was planned because we want to retro as much rubbish and redundant equipment as we can on the helicopter back to Kangerlussuaq.
Tom and Tris will be returning, not as rubbish and redundant field workers, but as valued and honoured field team members. Tom has been an invaluable field microbiologist and baker, and Tris has “got more spectra than he knows what to do with them”. It seems that the albedo monitoring and measuring has gone very well. Tris handed over the UAV controls to Joe yesterday, so the data acquisition by quadcopter will continue apace.
It’s been very busy for Joe and Johan too, who have been surface ice disk and ice core collecting these last two days. Johan reports that the DEM (digital elevation model) of the main (MODIS pixel-sized) sampling area was constructed yesterday too. Meanwhile, Jim is busy aerosol sampling. Microbiological incubations are ongoing at a pace.
The helicopter was out searching for a missing tourist yesterday, so our field crew change and dash to the snowline for sampling may be a little delayed today. We’re all hoping that the person will be found safely, and we can live with a delay if necessary. More on the dash next post with luck.